# Forecasting Galaxy Cluster H i Mass Recovery with CHIME at Redshifts z=1 and 2 via the IllustrisTNG Simulations

### v1

@@ Abstract
The Canadian Hydrogen Intensity Mapping Experiment (CHIME) is a drift-scan interferometer designed to map the entire northern sky every 24 hours. The all-sky coverage and sensitivity to neutral hydrogen flux at intermediate redshifts makes the instrument a resource for other exciting science in addition to cosmology for which it was originally designed. Characterizing the contents of CHIME’s beam-smoothed maps will aid in planning novel use-cases for the instrument, particularly those pertaining to galaxy evolution studies. Here, we demonstrate its utility for the study of the H i content of stacked galaxy populations across environments. Focusing first on galaxy clusters, we use simulated data from the IllustrisTNG project to understand the relative contribution of objects that fall into CHIME’s synthesized beam to the observed H i flux using stacking analyses at a couple of representative redshifts. We find that there is an appreciable difference in the estimated stacked flux when galaxy clusters or cluster member galaxies are used as tracers compared to stacking on a general galaxy catalog. Stacking on galaxy clusters, we report that it is possible to recover an average $M_{\mathrm{HI}}$ for clusters as a function of redshift and selection criteria.
@@ 1 Introduction
The Canadian Hydrogen Intensity Mapping Experiment (CHIME; CHIME Collaboration et al. 2022a) is a cylindrical drift-scan interferometer operating between 400-800 MHz at the Dominion Radio Astrophysical Observatory in Penticon, Canada. CHIME was designed to use the redshifted 21 cm  emission of neutral hydrogen to measure large scale structure via a low spatial resolution, high spectral resolution radio galaxy survey (referred to as intensity mapping; Battye et al. 2004; Chang et al. 2008; Wyithe & Loeb 2008; Peterson et al. 2009; Shaw et al. 2015) at frequencies corresponding to $0.82.5>z>$ 0.8. To this end, maps from CHIME cover $\sim$60 per cent of the Northern sky daily with sub-degree angular resolution ($0.52^{\circ}-0.26^{\circ}$ for the synthesized beam), which translates to physical scales of $50-10$ Mpc (Newburgh et al., 2014; CHIME Collaboration et al., 2022a).
@@
CHIME’s redshift band and sensitivity to neutral hydrogen could provide a new probe of the integrated H i content of galaxy clusters spanning the earliest assembling clusters around $z\sim 2$ to the quenching of star formation around $z\sim$1. Although ongoing and planned experiments are aimed at studying the evolution of cluster H i in individual galaxy clusters out to $z\sim 1.5$ (Ai et al., 2017), data from CHIME could be used to better understand the reservoir of neutral gas in dense environments and potentially to statistically distinguish the role of AGN feedback and whether it follows model predictions as a function of redshift (e.g., Villaescusa-Navarro et al., 2016; Tremmel et al., 2019; Dong et al., 2023).
@@
In preparation for cross-correlating CHIME data with galaxy cluster (or other) catalogs to answer questions related to galaxy evolution, or, more specifically, the evolution of the gaseous medium in and around galaxies, it is important to understand which class of objects will actually contribute to the measured H i flux in stacked data. Our low spatial resolution ensures that we are able to measure the integrated flux from objects that fall in the beam with the caveat that we are not able to resolve individual objects and/or their immediate environments. Instead, our target objects and surrounding structures are all likely to be included in the beam and in our stacking analyses. It is essential that we understand how much surrounding ancillary objects contaminate our stacked flux measurements. In this paper, we use the Illustris TNG300 suite of simulations of large scale structure and its neutral hydrogen content to estimate whether CHIME could be used to measure the H i content in galaxy clustesr across this interesting epoch. Despite CHIME’s low spatial resolution, which integrates cluster and field galaxies together, we find that it is possible to infer the average H i mass of stacked clusters and recover the relative mass ratio of galaxy clusters at different redshifts and/or of different cluster populations.
We assume the general cosmological parameters from Planck Collaboration et al. (2016) used by the TNG300 simulation (Pillepich et al., 2018; Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) of the IllustrisTNG project (Nelson et al., 2019), $h=0.6774$, $\Omega_{M}=0.3089$, and $\Omega_{\Lambda}=0.6911$, and we adopt the H iH2 transition prescription from Gnedin & Kravtsov (2011, see Section 2.1) to assign H i masses to each halo.
The structure of this paper is as follows: in Section 2, we discuss the simulated data that facilitates this analysis; in Section 3, we go through the steps of our stacking analysis; in Section 4, we lay out a means of recovering $M_{\mathrm{HI}}$ from CHIME galaxy cluster stacks; and in Section 5, we consider the broad applicability of this work.
@@ 2 Simulating CHIME observations
In this section we describe how we generate a mock CHIME observation of an IllustrisTNG simulation field. This paper is meant to describe a proof-of-concept reconstruction, and so we do not follow the simulation template procedure outlined in CHIME Collaboration et al. (2023), which was used to find parameter errors on model parameters such as the 21 cm power spectrum ($P_{\mathrm{HI},g}$), the bias of galaxy and 21 cm fields ($b_{g}$, $b_{\mathrm{HI}}$), etc., and included effects such as fingers-of-god, nor do we include processing steps such as filtering and weighting. Instead, we will adopt a less rigorous approach of convolving projections of the TNG300 volume with the synthesized CHIME beam shape and using that as a proxy for the stacking measurement to understand whether CHIME can in principle recover H i information from high redshift clusters.
@@ 2.1 TNG300 simulation from the IllustrisTNG project
We select the TNG300 simulation, part of the Illustris suite of cosmological simulations (Pillepich et al., 2018; Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) as the parent simulation for generating our mock CHIME maps due its large volume, native inclusion of H i , and relevant redshift coverage. In the TNG300 simulation, the evolution of the gas content - both molecular and atomic hydrogen - is tracked for every sufficiently massive subhalo across environments.
The TNG300 Molecular and atomic hydrogen (HI+H2) galaxy contents catalogs (Diemer et al., 2018, 2019) provide the neutral hydrogen mass of each subhalo at two discrete redshifts, $z=1$ and $z=2$, that also overlap with the CHIME observation band. Since $z=1.5$ (568 MHz) falls into the radio frequency interference-afflicted $500-585$ MHz sub-band and is inaccessible to us in analysis of real CHIME data, we do not include it in our mock data analysis either.
Figure 1: The relationship between $M_{\mathrm{star}}$ and $M_{\mathrm{HI}}$ for massive galaxies in both the simulated TNG300 sample and the observed ALFALFA-SDSS galaxy catalog at $z=0$.
In order to ensure that the H i masses in the Molecular and atomic hydrogen (HI + H2) galaxy contents catalogs are in line with observations, we compare the $M_{\mathrm{star}}-M_{\mathrm{HI}}$ relation from TNG300 to the derived masses in the ALFALFA-SDSS Galaxy Catalog (Durbala et al., 2020). We use the $z=0$ TNG300 hydrogen gas mass catalog here to match the low redshift coverage of H i in ALFALFA ($z\leq 0.06$; Giovanelli et al. 2005).
Looking at the galaxies with $M_{\mathrm{star}}\geq 10^{10}\,M_{\odot}$ that we are including in our analyses here, we see that the TNG300 galaxies are qualitatively consistent with the stellar mass-H i mass observed in ALFALFA-SDSS galaxies (Fig. 1). The discrepancy at low neutral hydrogen mass is likely due to ALFALFA-SDSS galaxies originating in an H i survey with limited sensitivity (down to $M_{\mathrm{HI}}\sim 10^{8}\,M_{\odot}$ at 20 Mpc), while the TNG300 H i model we adopt (Gnedin & Kravtsov, 2011, alternately referred to as GK11) has no such constraints. The similarity of TNG300 H i abundances to observations is discussed further in Diemer et al. (2019).
@@ 2.2 Constructing a beam-like convolutional kernel
To generate the low-resolution CHIME intensity maps from high-resolution TNG300 simulations, we require a frequency-dependent synthesized beam model of CHIME. The cylindrical design of the CHIME reflector gives different cross-sections in the North-South (N-S) and East-West (E-W) directions.
@@
We use the form of the synthesized E-W beam model from CHIME Collaboration et al. (2023) at 710 MHz ($z=1$), with a few simplifications: we have limited the beam to hour angles less than ±$\pm 4^{\circ}$ which neglects the effects of aliasing. This beam is computed by beamforming to declination $\delta=0^{\circ}$, and we use only the Y-polarized beam and include only inter-cylinder baselines (baselines that span more than one cylinder). The full form of the synthesized beam is described in more detail in Section 4.3.4 of CHIME Collaboration et al. (2023). For the E-W beam at 473 MHz ($z=2$), we use full width at half maximum (FWHM) $\propto 1/\nu$ to re-scale the 710 MHz ($z=1$) beam model. In this work we interpolate the beam function to a coarser grid-size to convolve more easily with the simulations.
The two-dimensional beam is then the outer product of the N-S and E-W cross-sections. Although this does not capture the rich structure of the CHIME beam (CHIME Collaboration et al., 2022b), the simplified beam is sufficient for the explorations in this paper. These synthesized (and cross-sectional) beam models are shown in Fig. 2.
@@ 2.3 Generating mock intensity maps
Here we outline the procedure adopted to generate mock intensity maps from the simulated data. Assuming that the H i gas is optically thin in the sources we are examining, we can use the formalism from Wild (1952) to derive the neutral hydrogen intensity from H i mass (Draine, 2011):
@@
The luminosity distance, $D_{L}$, is defined consistent with our chosen cosmology and the simulation volume – for $z=1$, $D_{L}$ = $6.8\times 10^{3}$ to $7.0\times 10^{3}$ Mpc and for $z=2$, $D_{L}$ = $1.59\times 10^{4}$ to $1.60\times 10^{4}$ Mpc. We convert to flux density in Jy by dividing off the width of the CHIME frequency sub-band (0.390 MHz), which assumes that we know the spectral position of targets to within $dz\lesssim 0.001$ at $z=1$ and $dz\lesssim 0.002$ at $z=2$. Selecting a different range in redshift/frequency will impact the noise properties of the stacks as well as the observed flux. Foreground filtering strongly attenuates the signal with increasing bandwidth. To approximate this effect in our mock observations, we apply a corrective factor of 0.2 (CHIME Collaboration et al., 2023) to the flux. This is for our narrowest CHIME band, comparable to the precision of a spectroscopic catalog. More advanced foreground filtering may allow for a greater tolerance in $dz$ in the future.
[…]
To quantify the contribution of different classes of objects to the measured neutral hydrogen flux, we generate population-specific TNG300 catalogs – i.e., only field galaxies or only cluster member galaxies.
Given the limited angular extent of each TNG300 slice at $z=1$ and $z=2$ we tile rotated (and flipped) versions of the same projected map to create a larger map to smooth and stack. Since we also permute the projected maps from the existing (300 Mpc)3 simulation volume with respect to the reference position used in calculating $F_{\mathrm{obs}}$, the final tiled map encompasses 384 times the area of a single projected simulation slice, totalling $\sim 4100$ deg2 at $z=1$ and $\sim 1900$ deg2 at $z=2$. (See Appendix A for more details.) It is this larger intensity map that is then convolved with the CHIME beam model. The mock sky coverage is comparable to the extent of the eBOSS fields used in CHIME Collaboration et al. (2023). We add Gaussian noise with $\sigma=0.6$ mJy (CHIME Collaboration et al., 2023) to the beam-convolved map in order to better replicate CHIME observations.
Fig. 3 shows the progression from the higher resolution native TNG300 mock intensity map to the beam-smoothed mock intensity map with realistic noise. We note that we do not include foreground emission or filtering, frequency masking from radio frequency interference, or point source removal. Instead, the results we present provide a proof-of-principle extraction of astrophysical neutral hydrogen in clusters from intensity mapping experiments like CHIME.
@@ 3 Mock stacking analyses
We stack the signal from the simulated H i maps for the two redshifts separately. For each source in the simulation, we create a two-dimensional, $6^{\circ}\times 6^{\circ}$ cutout in RA and Dec, following Section 4.7 of CHIME Collaboration et al. (2023). These cutouts are then directly co-added, returning a ‘stacked’ signal from the average of the co-added signals (see Fig. 4 for an example). The intensity map area of our $z=1$ mock is comparable to the area used in the eBOSS/LRG and QSO cross-matched stacks from CHIME Collaboration et al. (2023), while the area of our $z=2$ mock is comparable to the area used in the eBOSS/ELG cross-matched stack. We do not expect that this difference in area impacts our ability to compare mock stacking analyses across redshifts, given that the signal we are measuring is innately normalized by virtue of averaging the stacks.
@@
    radius-selected clusters; stacked on radius-selected galaxy clusters with haloes with comoving $r_{\mathrm{vir}}>0.75\,h^{-1}$ Mpc, independently determined for $z=1$ and $z=2$.
@@
We show example stacks in Fig. 4. The top row includes stacks (i) - (iii) and (vii) on the full catalog returned for the simulation volume, and the bottom row shows the same stacks for a more realistic number of randomly selected galaxies in the simulation volume. These realistic stacks are comparable to those in CHIME Collaboration et al. (2023). The amplitude of the stacked signal is comparable between the full catalog and realistic stacks, with the noise properties varying as predicted based on the number of targets used in each stack.

### v2

@@ Abstract
The Canadian Hydrogen Intensity Mapping Experiment (CHIME) is a drift-scan interferometer designed to map the entire northern sky every 24 hours. The all-sky coverage and sensitivity to neutral hydrogen flux at intermediate redshifts makes the instrument a resource for other exciting science in addition to cosmology for which it was originally designed. Here, we demonstrate its utility for the study of the H i content of galaxy populations across environments and redshifts. We use simulated data from the IllustrisTNG project to generate mock CHIME-like intensity maps, which we cross-correlate with various tracers – including galaxies and galaxy clusters – to recover aggregate H i signals from stacking analyses. We find that there is more flux in stacks on galaxy clusters or cluster member galaxies compared to those on a general galaxy catalog due to the enhanced number of H i-rich sources included in the CHIME primary beam. We report that it is possible to infer an average $M_{\mathrm{HI}}$ for clusters as a function of redshift and selection criteria from the signal in their averaged stacks despite the instrument’s low spatial resolution. This proof-of-concept result opens up a promising, and timely, new avenue to measure the evolution of the neutral hydrogen content in intermediate-to-high redshift galaxy clusters via cross-correlation of galaxy cluster catalogs with 21 cm  intensity maps.
@@ 1 Introduction
The Canadian Hydrogen Intensity Mapping Experiment (CHIME; CHIME Collaboration et al. 2022a) is a cylindrical drift-scan interferometer operating between 400-800 MHz at the Dominion Radio Astrophysical Observatory in Penticton, Canada. CHIME was designed to use the redshifted 21 cm  emission of neutral hydrogen to measure large scale structure via a low spatial resolution, high spectral resolution survey (referred to as intensity mapping; Battye et al. 2004; Chang et al. 2008; Wyithe & Loeb 2008; Peterson et al. 2009; Shaw et al. 2015) at frequencies corresponding to $0.82.5>z>$ 0.8. To this end, maps from CHIME cover $\sim$60 per cent of the Northern sky daily with sub-degree angular resolution ($0.52^{\circ}-0.26^{\circ}$ for the synthesized beam), which translates to physical scales of $50-10$ Mpc (Newburgh et al., 2014; CHIME Collaboration et al., 2022a).
@@
CHIME’s redshift band and sensitivity to neutral hydrogen could provide a new probe of the integrated H i content of galaxy clusters spanning the earliest assembling clusters around $z\sim 2$ to the quenching of star formation around $z\sim$1. At high redshift, galaxy clusters are generally thought to be gas-rich and contain large reservoirs of neutral hydrogen (e.g., Francis et al., 2001; Battye et al., 2004). There are ongoing and planned experiments aimed at studying the evolution of cluster H i in individual galaxy clusters. The Five-hundred-meter Aperture Spherical radio Telescope (FAST, Jiang et al., 2019, 2020; Qian et al., 2020) should be able to measure the H i signal from individual neutral hydrogen-rich galaxy clusters out to $z\sim 1.5$ (Ai et al., 2017). However, even the best-case projected integration times for FAST to reach the required depth are on the order of hours to days, depending upon cluster redshift, which limits potential survey scope. Data from CHIME can be used to better understand the average reservoir of neutral gas in dense environments at the population level. Such measurements potentially have the power to statistically distinguish the role of AGN feedback in driving the evolution of clusters’ neutral hydrogen reservoir and confirm whether it follows model predictions as a function of redshift (e.g., Villaescusa-Navarro et al., 2016; Tremmel et al., 2019; Dong et al., 2023).
@@
In preparation for cross-correlating CHIME data with galaxy cluster (or other) catalogs to answer questions related to galaxy evolution, or, more specifically, the evolution of the gaseous medium in and around galaxies, it is important to understand which objects actually contribute to the measured H i flux in stacked data. Our low spatial resolution ensures that we are able to measure the integrated flux from objects that fall in the beam; however, that low resolution comes with the caveat that we are not able to resolve individual objects and/or their immediate environments. Instead, structures surrounding our target objects are also likely to be included in the beam and in our stacking analyses. It is essential that we understand how much surrounding ancillary objects contaminate our stacked flux measurements. In this paper, we use the Illustris TNG300 suite of simulations of large scale structure and its neutral hydrogen content to estimate whether CHIME could be used to measure the H i content in galaxy clusters across this interesting epoch.
We assume the general cosmological parameters from Planck Collaboration et al. (2016) used by the TNG300 simulation (Pillepich et al., 2018; Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) of the IllustrisTNG project (Nelson et al., 2019), $h=0.6774$, $\Omega_{M}=0.3089$, and $\Omega_{\Lambda}=0.6911$, and we adopt the H iH2 transition prescription from Gnedin & Kravtsov (2011) to assign H i masses to each halo.
The structure of this paper is as follows: in Section 2, we discuss the simulated data that facilitates this analysis; in Section 3, we go through the steps of our stacking analysis; in Section 4, we lay out an example means of recovering $M_{\mathrm{HI}}$ from CHIME galaxy cluster stacks; and in Section 5, we consider the broad applicability of this work.
@@ 2 Simulating CHIME observations
In this section we describe how we generate a mock CHIME observation of an IllustrisTNG simulation field. This paper is meant to describe a proof-of-concept reconstruction, and so we do not follow the simulation template procedure outlined in CHIME Collaboration et al. (2023), which was used to find parameter errors on model parameters such as the 21 cm power spectrum ($P_{\mathrm{HI},g}$), the bias of galaxy and 21 cm fields ($b_{g}$, $b_{\mathrm{HI}}$), etc., and included effects such as fingers-of-god, nor do we include processing steps such as filtering and weighting. Instead, we adopt a less rigorous approach of convolving projections of the TNG300 volume with the synthesized CHIME beam shape and using that as a proxy for the stacking measurement to understand whether CHIME can in principle recover H i information from high redshift clusters.
@@ 2.1 TNG300 simulation from the IllustrisTNG project
We select the TNG300 simulation, part of the Illustris suite of cosmological simulations, as the parent simulation for generating our mock CHIME maps due its large volume, native inclusion of H i, and relevant redshift coverage. In the TNG300 simulation, the evolution of the gas content - both molecular and atomic hydrogen - is tracked for every sufficiently massive ($M_{\mathrm{star}}\geq 5\times 10^{10}M_{\odot}$ or $M_{\mathrm{gas}}\geq 10^{9}M_{\odot}$) subhalo across environments.
The TNG300 Molecular and atomic hydrogen (HI+H2) galaxy contents catalogs (Diemer et al., 2018, 2019) provide the neutral hydrogen mass of each subhalo at two discrete redshifts, $z=1$ and $z=2$, that also overlap with the CHIME observation band. Since $z=1.5$ (568 MHz) falls into the radio frequency interference-afflicted $500-585$ MHz sub-band and is inaccessible to us in analysis of real CHIME data, we do not include it in our mock data analysis either. We select the Gnedin & Kravtsov (2011, alternately referred to as GK11) H i model from those included in the Molecular and atomic hydrogen (HI+H2) galaxy contents catalogs; the similarity of TNG300 H i abundances to observations is discussed further in Diemer et al. (2019).
@@ 2.2 Constructing a beam-like convolutional kernel
To generate the low-resolution CHIME intensity maps from high-resolution TNG300 simulations, we require a frequency-dependent synthesized beam model of CHIME. The cylindrical design of the CHIME reflector has a different cross-sectional beam profile in the North-South (N-S) and East-West (E-W) directions.
@@
We use the synthesized E-W beam model from CHIME Collaboration et al. (2023) at 710 MHz ($z=1$), with a few simplifications: we have limited the beam to hour angles less than ±$\pm 4^{\circ}$ which neglects the effects of aliasing; this beam is computed by beamforming to declination $\delta=0^{\circ}$, and we use only the Y-polarized beam and include only inter-cylinder baselines (baselines that span more than one cylinder). The full form of the synthesized beam is described in more detail in Section 4.3.4 of CHIME Collaboration et al. (2023). For the E-W beam at 473 MHz ($z=2$), we use full width at half maximum (FWHM) $\propto 1/\nu$ to re-scale the 710 MHz ($z=1$) beam model. We also interpolate the beam function to a coarser grid-size to convolve more easily with the simulations.
The two-dimensional, synthesized beam is then the outer product of the N-S and E-W cross-sections. Although this does not capture the rich structure of the CHIME beam (CHIME Collaboration et al., 2022b), the simplified beam is sufficient for the explorations in this paper as we focus here on the polarization-independent flux measured in the primary component of the synthesized beam. These beam models, both synthesized and cross-sectional, are shown in Fig. 1.
@@ 2.3 Generating mock intensity maps
Here we outline the procedure adopted to generate mock intensity maps from the simulated data. Assuming that the H i gas is optically thin in the sources we are examining, we use the formalism from Wild (1952) to derive the neutral hydrogen intensity from H i mass (Draine, 2011):
@@
The observed flux, $F_{\mathrm{obs}}$, is defined as the flux density integrated over the frequency range of the relevant CHIME sub-band. This flux should be equivalent to Stokes I, which combines the two linear polarizations for total intensity ($I=X^{2}+Y^{2}$); for the purposes of this paper, we use the CHIME measured Y polarization beam as a proxy for the Stokes I beam due to its well-understood behavior in these two redshift regimes, noting that the two instrumental polarizations are shown to return consistent results in CHIME Collaboration et al. (2023), so that any difference in the X- and Y-polarized beams should have a negligible impact on the analysis in this work.
The luminosity distance, $d_{L}$, is defined consistent with our chosen cosmology and the simulation volume – for $z=1$, $d_{L}$ = $6.8\times 10^{3}$ to $7.0\times 10^{3}$ Mpc and for $z=2$, $d_{L}$ = $1.59\times 10^{4}$ to $1.60\times 10^{4}$ Mpc. We convert to flux density in Jy by dividing off the width of the CHIME frequency sub-band (0.390 MHz), which assumes that we know the spectral position of targets to within $dz\lesssim 0.001$ at $z=1$ and $dz\lesssim 0.002$ at $z=2$. In stacking, see Section 3, we assume that all objects in the mock intensity map are at $z=1\pm dz$ or 2 ±$\pm dz$. Selecting a different range in redshift/frequency will impact the noise properties of the stacks as well as the observed flux. Foreground filtering strongly attenuates the signal with increasing bandwidth. To approximate this effect in our mock observations, we apply a corrective factor of 0.2 to the flux, corresponding to the mean foreground filtering attenuation between catalogs used in CHIME Collaboration et al. (2023). This is for our narrowest CHIME band, comparable to the precision of a spectroscopic catalog. More advanced foreground filtering may allow for a greater tolerance in $dz$ in the future.
@@
Given the limited angular extent of each TNG300 slice at $z=1$ and $z=2$, we tile rotated (and flipped) versions of the same projected map to create a larger map to smooth and stack. Since we also permute the projected maps from the existing (300 Mpc)3 simulation volume with respect to the reference position used in calculating $F_{\mathrm{obs}}$, the final tiled map encompasses 384 times the area of a single projected simulation slice, totalling $\sim 4100$ deg2 at $z=1$ and $\sim 1900$ deg2 at $z=2$. (See Appendix A for more details.) It is this larger intensity map that is then convolved with the CHIME beam model. The mock sky coverage is comparable to the extent of the eBOSS fields used in CHIME Collaboration et al. (2023). We add Gaussian noise with $\sigma=0.6$ mJy (CHIME Collaboration et al., 2023) to the beam-convolved map in order to better replicate CHIME observations.
Fig. 2 shows the progression from the higher resolution native TNG300 mock intensity map to the beam-smoothed mock intensity map with realistic noise. We note that we do not include foreground emission, frequency masking from radio frequency interference, or point source removal. Instead, the results we present provide a proof-of-principle extraction of astrophysical neutral hydrogen in clusters from intensity mapping experiments like CHIME.
@@ 3 Mock stacking analyses
We stack the signal from the simulated H i maps for the two redshifts separately. For each source in the simulation, we create a two-dimensional, $6^{\circ}\times 6^{\circ}$ cutout in RA and Dec, following Section 4.7 of CHIME Collaboration et al. (2023). These cutouts are then directly co-added, returning a ‘stacked’ signal from the average of the co-added signals (see Fig. 3 for an example). The intensity map area of our $z=1$ mock is comparable to the area used in the eBOSS/LRG and QSO cross-matched stacks from CHIME Collaboration et al. (2023), while the area of our $z=2$ mock is comparable to the area used in the eBOSS/ELG cross-matched stack. We do not expect that the $z$-dependent difference in area impacts our ability to compare mock stacking analyses across redshifts, given that the signal we are measuring is innately normalized by virtue of averaging the stacks.
@@
    radius-selected clusters; stacked on radius-selected galaxy clusters with haloes with comoving $R_{200}>0.75\,h^{-1}$ Mpc, independently determined for $z=1$ and $z=2$.
@@
We show example stacks in Fig. 3. The top row includes stacks (i) - (iii) and (vii) on the full catalog returned for the simulation volume, and the bottom row shows the same stacks for a more realistic number of randomly selected galaxies in the simulation volume. These realistic stacks are directly comparable to those in CHIME Collaboration et al. (2023), including the same number of target galaxies. The amplitude of the stacked signal is equivalent between the full catalog and realistic stacks, with the noise properties varying as predicted based on the number of targets used in each stack. The implications of stacking on unresolved target source is described in Section 4.